A company prepares two types of packed salads, A and B. During any week, it can prepare at most 650 salads. In week 1 the cost per salad A is S4 and its demand is 400, while the cost per salad B is S3.50 and its demand is 200. In week 2 the cost per salad A is $4.40 and its demand is 300, while the cost per salad B is $4 and its demand is 300. The demands must be met on time. Also, it costs 50e to hold a salad A, and 40e to hold a salad B, in inventory for a week. Formulate an LP to minimize the total cost of meeting the next two weeks demands
Answer #1

Let Tla and Tib be the number of products of A and B produced in the week 1, respectively

Let x_{2a} \,\,\,and\,\,\, x_{2b} be the number of products of A and B produced in the week 1, respectively.

Let I_a \,\,\,\&\,\,\,I_b , be the number of products of A and B produced in the week 1 and kept as inventory for next week respectively.

Objective function:

Minimize: Total cost Z = 4x_{1a}+3.50x_{1b}+4.4x_{2a}+4x_{2b}+0.5I_{1a}+0.4I_{b}

Subject to:

  1. Capacity constraints:

x_{1a}+x_{1b} \leq 650

x_{2a}+x_{2b} \leq 650

2. Demand Constraints:

x_{1a} \geq 400

x_{1b} \geq 200

x_{2a} +I_a \geq 300

x_{2b} +I_b \geq 300

I_a = x_{1a} -400

I_b = x_{1b} -200

I_a,I_b, x_{1a}, x_{1b}, x_{2a} \,\,\, and \,\,\, x_{2b} \geq 0

NOTE: Above model can also be modelled as below:

Objective function:

Minimize: Total cost Z = 4x_{1a}+3.50x_{1b}+4.4x_{2a}+4x_{2b}+0.5(x_{1a}-400)+0.4(x_{1b} -200)

Subject to:

  1. Capacity constraints:

x_{1a}+x_{1b} \leq 650

x_{2a}+x_{2b} \leq 650

2. Demand Constraints:

x_{1a} \geq 400

x_{1b} \geq 200

x_{2a} +x_{1a}-400 \geq 300 or x_{2a} +x_{1a} \geq 700

x_{2b} +x_{1b}-200 \geq 300 or x_{2b} +x_{1b} \geq 500

x_{1a}, x_{1b}, x_{2a} \,\,\, and \,\,\, x_{2b} \geq 0